# Detecting Out-of-distribution through the Lens of Neural Collapse

### v4

@@ 4.3 Ablation on the Filtering Effect
In Table 3, we assess different orders of $p$-norm as the filtering term, compared to the $L1$ norm used so far. To ensure a fair comparison, we report the best performance from the filter strengths $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. The rest of the setup follows the ImageNet benchmarks in Section 4.1. As shown in Table  3, filtering with $L1$ norm achieves the best performance across OOD datasets, aligning with prior observations Huang et al. (2021); Park et al. (2023). Meanwhile, we observe that in rare scenarios, e.g., a ResNet-18 on CIFAR-10, the $\mathtt{L1}$ norm cannot effectively characterize OOD’s proximity to the origin, leading to no extra performance gain compared to simply thresholding on $\mathtt{pScore}$. In these cases, our algorithm benefits from its ability to automatically select a low filter strength based on validation results, effectively disregarding the filtering term.
@@ 5 Related Work
OOD Detection Extensive research has focused on developing OOD detection algorithms. One line of work is post-hoc and builds upon pre-trained models. For example, Hendrycks et al. (2019); Liang et al. (2018); Liu et al. (2020); Sun et al. (2021); Sun and Li (2022) design OOD score over the output space of a classifier. Meanwhile, Lee et al. (2018) and Sun et al. (2022) measure OOD-ness from the perspective of ID clustering in feature space. Our work extends the observation that ID features tend to cluster from the perspective of Neural Collapse. While existing work is more focused are certain classification tasks than others, our proposed OOD detector is tested to be highly versatile.
Another line of work explores the regularization of OOD detection in training. For example, DeVries and Taylor (2018); Hsu et al. (2020) propose OOD-specific architecture whereas Huang and Li (2021); Wei et al. (2022) design OOD-specific training loss. In particular, Tack et al. (2020) brings attention to representation learning for OOD detection and proposes an OOD-specific contrastive learning scheme. Our work does not belong to this school of thought and is not restricted to specific training schemes or architecture.
Neural Collapse Neural Collapse was first observed in Papyan et al. (2020). During Neural Collapse, the penultimate layer features collapse to class means, the class means and the classifier collapses to a simplex equiangular tight framework, and the classifier simplifies to adopt the nearest class-mean decision rule. Further work provides theoretical justification for the emergence of Neural Collapse (Han et al., 2021; Mixon et al., 2020; Zhou et al., 2022; Zhu et al., 2021). In addition, Zhu et al. (2021) derives an efficient training algorithm drawing inspiration from Neural Collapse. Our concurrent work Ammar et al. (2023) also leverages insights from Neural Collapse for OOD detection. However, they tackle from the subspace perspective and largely overlook class-specific information revealed by Neural Collapse, which is essential for our work.
@@ 6 Conclusion
Despite significantly reducing the generalization discrepancy, our $\mathtt{NCI}$ does not completely close the gap. For example, there is still an absolute difference of 1.72% in AUROC compared to the best result on the ImageNet ResNet-50 benchmark.
@@ A.3 ImageNet
Swin v2 In Table 2, we use the $\mathtt{timm}$ Wightman (2019) implementation of Swin v2 as well as their pre-trained checkpoint ’swinv2_base_window8_256’.
@@ Appendix E The Prevalence of Neural Collapse across Canonical Classification Tasks
The phenomenon of Neural Collapse, as established in the seminal work by Papyan et al. Papyan et al. (2020) and corroborated by subsequent studies Han et al. (2021); Mixon et al. (2020); Zhou et al. (2022); Zhu et al. (2021), widely exists across canonical classification datasets and model architectures. The prevalent occurrence of Neural Collapse forms a robust foundation for the design of our versatile OOD detectors. To this end, we review the empirical evidence of Neural Collapse across different datasets and model architectures in Figure 3, Figure 4, Figure 5, Figure 6, and Figure 7. Comparing CIFAR-10 and ImageNet behaviors with ResNet backbone in Figure 7, we note that the clustering of CIFAR-10 is more prominent than Imagenet, as indicated by a higher ratio of between-class variance to within-class covariance. Note that the figures and captions are sourced from Papyan et al. (2020). The definition and notation follow Section 3.

### v5

@@ 4.3 Ablation on the Filtering Effect
In Table 3, we assess different orders of $p$-norm as the filtering term, compared to the $L1$ norm used so far. To ensure a fair comparison, we report the best performance from the filter strengths $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. The rest of the setup follows the ImageNet benchmarks in Section 4.1. As shown in Table  3, filtering with $L1$ norm achieves the best performance across OOD datasets, aligning with prior observations (Huang et al., 2021; Park et al., 2023). Meanwhile, we observe that in rare scenarios, e.g., a ResNet-18 on CIFAR-10, the $\mathtt{L1}$ norm cannot effectively characterize OOD’s proximity to the origin, leading to no extra performance gain compared to simply thresholding on $\mathtt{pScore}$. In these cases, our algorithm benefits from its ability to automatically select a low filter strength based on validation results, effectively disregarding the filtering term.
@@ 5 Related Work
OOD Detection Extensive research has focused on developing OOD detection algorithms. One line of work is post-hoc and builds upon pre-trained models. For example, Hendrycks et al. (2019); Liang et al. (2018); Liu et al. (2020); Sun et al. (2021), and Sun and Li (2022) design OOD score over the output space of a classifier. Meanwhile, Lee et al. (2018) and Sun et al. (2022) measure OOD-ness from the perspective of ID clustering in feature space. Our work extends the observation that ID features tend to cluster from the perspective of Neural Collapse. While existing work is more focused are certain classification tasks than others, our proposed OOD detector is tested to be highly versatile.
Another line of work explores the regularization of OOD detection in training. For example, DeVries and Taylor (2018) and Hsu et al. (2020) propose OOD-specific architecture whereas Huang and Li (2021) and Wei et al. (2022) design OOD-specific training loss. In particular, Tack et al. (2020) brings attention to representation learning for OOD detection and proposes an OOD-specific contrastive learning scheme. Our work does not belong to this school of thought and is not restricted to specific training schemes or architecture.
Neural Collapse Neural Collapse was first observed in Papyan et al. (2020). During Neural Collapse, the penultimate layer features collapse to class means, the class means and the classifier collapses to a simplex equiangular tight framework, and the classifier simplifies to adopt the nearest class-mean decision rule. Further work provides theoretical justification for the emergence of Neural Collapse (Han et al., 2021; Mixon et al., 2020; Zhou et al., 2022; Zhu et al., 2021). In addition, Zhu et al. (2021) derives an efficient training algorithm drawing inspiration from Neural Collapse. Our concurrent work, Ammar et al. (2023), also leverages insights from Neural Collapse for OOD detection. However, they tackle from the subspace perspective and largely overlook class-specific information revealed by Neural Collapse, which is essential for our work.
@@ 6 Conclusion
Despite significantly reducing the generalization discrepancy, our $\mathtt{NCI}$ does not completely close the gap. Particularly, there is still an absolute difference of 1.72% in AUROC compared to the best result on the ImageNet ResNet-50 benchmark.
@@ A.3 ImageNet
Swin v2 In Table 2, we use the $\mathtt{timm}$ (Wightman, 2019) implementation of Swin v2 as well as their pre-trained checkpoint ’swinv2_base_window8_256’.
@@ Appendix E The Prevalence of Neural Collapse across Canonical Classification Tasks
The phenomenon of Neural Collapse, as established in the seminal work by Papyan et al. (2020) and corroborated by subsequent studies (Han et al., 2021; Mixon et al., 2020; Zhou et al., 2022; Zhu et al., 2021), widely exists across canonical classification datasets and model architectures. The prevalent occurrence of Neural Collapse forms a robust foundation for the design of our versatile OOD detectors. To this end, we review the empirical evidence of Neural Collapse across different datasets and model architectures in Figure 3, Figure 4, Figure 5, Figure 6, and Figure 7. Comparing CIFAR-10 and ImageNet behaviors with ResNet backbone in Figure 7, we note that the clustering of CIFAR-10 is more prominent than Imagenet, as indicated by a higher ratio of between-class variance to within-class covariance. Note that the figures and captions are sourced from Papyan et al. (2020). The definition and notation follow Section 3.